pooja195 (pooja195):

Ok one thing im confused on is scientific notation how do i know if the exponent will be negative or positive?If this makes sense at all ;-;

OpenStudy (misty1212):

\[10^2=100\\ 10^{-2}=0.01\]

OpenStudy (misty1212):

so \[200=2\times 10^2\] and \[0.02=2\times 10^{-2}\]

OpenStudy (acxbox22):

if the number itself is less than 1 than the power is negative and vice versa

pooja195 (pooja195):

ah ok that makes sense thanks :) was just a bit confused on this thanks for clearing it up guys!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\] (i guess i am a guy today)